# Sample size and power calculations for causal inference in observational studies

### v3

@@ Abstract
This paper investigates the theoretical foundation and develops analytical formulas for sample size and power calculations for causal inference with observational data. By analyzing the variance of the inverse probability weighting estimator of the average treatment effect, we decompose the power calculations into three components: propensity score distribution, potential outcome distribution, and their correlation. We show that to determine the minimal sample size of an observational study, it is sufficient under mild conditions to have two parameters additional to the standard inputs in the power calculation of randomized trials, which quantify the strength of the confounder-treatment and the confounder-outcome association, respectively. For the former, we propose using the Bhattacharyya coefficient, which measures the covariate overlap and, together with the treatment proportion, leads to a uniquely identifiable and easily computable propensity score distribution. For the latter, we propose a sensitivity parameter bounded by the R-squared statistic of the regression of the outcome on covariates. Utilizing the Lyapunov Central Limit Theorem on the linear combination of covariates, our procedure does not require distributional assumptions on the multivariate covariates. We develop an associated R package PSpower.
@@ 1 Introduction
Randomized controlled trials are the gold standard for causal inference. Designing randomized trials requires sample size and power calculations, in which investigators pre-specify a target estimand—typically the effect size of a treatment—and characteristics of the sample of the intended study. Then the goal is to determine the minimal sample size needed to achieve the statistical power ($1-\beta$), given a type I error rate ($\alpha$) level, in testing the treatment effect estimated by a specific method based on a sample that satisfies these characteristics, or conversely, to determine the power in detecting a treatment effect at the $\alpha$ level given a fixed sample size. Sample size and power calculations are the two sides of the same coin and we will use them interchangeably henceforth. Power calculation is conducted in the design stage of a study, prior to data collection. Thus, individual data are usually unavailable; investigators at best have some summary information of previous similar studies, e.g. mean and variances of covariates and outcomes by treatment arm. Analysts must rely on limited summary information of the intended study to recover the full data structure. The power calculation for randomized trials with treatment at one time has been well established and is routinely performed (Chow et al.,, 2017; Branson et al.,, 2024).
When randomized trials are not feasible due to logistical or ethical constraints, researchers are increasingly using observational data to emulate randomized trials. The power analysis for observational studies has been less developed in the literature. In fact, its role and utility has been questioned by some authors (Hernán,, 2022), who argued that the primary consideration in observational studies is to reduce confounding bias rather than improve efficiency or power. However, we view power analysis as crucial in designing prospective observational studies, which are usually subject to cost and time constraints. For example, in studies of rare diseases, recruiting patients is time-consuming, and thus investigators must determine the minimal sample size to achieve a reliable estimate before conducting the study. Another example concerns commercial population health data (such as the Truveta electronic health record data or IBM’s MarketScan health insurance claims data), which usually charge users cost proportional to the data size. With limited resources, a sample size calculation is critical for designing observational studies based on such data sources. Power analysis also helps researchers to understand the scope of data required and guide database selection in the design stage.
@@
Power analysis for causal inference with observational studies is challenging because of confounding. In randomized trials, there is no confounding by design; therefore, one can simply use the difference in the mean outcomes between arms to estimate the treatment effects. Closed-form formulas, e.g., based on the two-sample z-test, are readily applicable for sample size calculation. In contrast, in observational studies, confounding must be adjusted to obtain an unbiased estimate of treatment effects. Applying power calculation formulas designed for randomized trials to observational studies often leads to severely underpowered studies. In practice, a general power analysis method is to use simulations to generate individual data based on pre-specified summary statistics (Qin,, 2024). This method often makes unverifiable distributional assumptions of the multivariate covariates and lacks theoretical basis on what summary statistics are adequate to simulate the entire data. In fact, many different data generating models may satisfy the same summary statistics but lead to vastly different sample sizes; choosing between these models is usually ad hoc. Another method is to use the variance inflation factor (Hsieh and Lavori,, 2000; Scosyrev and Glimm,, 2019; Shook-Sa and Hudgens,, 2022; Austin,, 2021) to approximate the potential inflation in the sample size of an observational study compared to a randomized trial. In particular, Shook-Sa and Hudgens, (2022) derived a rigorous variance inflation factor formula based on the analysis of the variance of the inverse probability weighting estimator of the average treatment effect; their formula requires the distribution of the weights, which is often not available.
@@
In summary, the existing power calculation methods usually require information on the data generating process or the distributions of the weights or covariates, which is not always available in the design stage of a study. In this paper, we bypass this limitation by–under mild conditions–reducing the variance formula to a function of a few interpretable and commonly available summary parameters instead of the weights or covariates. Based on the analysis of variance of the inverse probability weighting estimator, we decompose power calculations into three components: propensity score distribution, potential outcome distribution, and their correlation. We show that, to determine the minimal sample size of an emulated trial with treatment at a single time, in addition to the standard inputs in the power calculation of a randomized trial, it is sufficient to have two parameters, which quantify the strength of the confounder-treatment and the confounder-outcome association, respectively. For the former, we propose to use the Bhattacharyya coefficient, which measures the covariate overlap and, together with the treatment proportion, leads to a uniquely identifiable and easily computable propensity score distribution. For the latter, we propose a sensitivity parameter bounded by the R-squared statistic of the regression of the outcome on covariates. We appeal to the Lyapunov Central Limit Theorem to reduce the multivariate covariates into a single Gaussian variable, which allows our procedure not to require specific distributional assumptions of the covariates. We show that a special case of our formula is the formula by Shook-Sa and Hudgens, (2022), which formalized the confounder-treatment association but not the confounder-outcome association. We develop an associated R package PSpower, available at https://CRAN.R-project.org/package=PSpower.
@@ 2 Variance and sample size based on the Hájek estimator
Suppose we have a sample of $N$ units. Each unit $i$ ($i=1,2,\ldots,N$) has a binary treatment indicator $Z_{i}$, with $Z_{i}=0$ being the control and $Z_{i}=1$ being treated, and a vector of $p$ covariates $X_{i}=(X_{1i},\cdots,X_{pi})^{\prime}$. For each unit $i$, there are two potential outcomes $\{Y_{i}(1),Y_{i}(0)\}$ mapped to treatment and control status, of which only the one corresponding to the observed treatment is observed, denoted by $Y_{i}=Z_{i}Y_{i}(1)+(1-Z_{i})Y_{i}(0)$; the other potential outcome is missing. The propensity score is the probability of a unit being assigned to the treatment group given the covariates (Rosenbaum and Rubin,, 1983): $e(x)=\Pr(Z_{i}=1\mid X_{i}=x)$.
The goal is to use the sample to estimate the average treatment effect:
@@
where $\tau(x)=\operatorname{\mathbb{E}}\{Y_{i}(1)-Y_{i}(0)\mid X_{i}=x\}$ is the conditional average treatment effect at covariate value $x$. To identify $\tau^{\mbox{\tiny{ATE}}}$, we maintain the two standard assumptions: (i) unconfoundededness, $\Pr(Z_{i}\mid Y_{i}(0),Y_{i}(1),X_{i})=\Pr(Z_{i}\mid X_{i})$; (ii) overlap, i.e. $10<\Pr(Z_{i}\mid X_{i})<$ 1. Then we can identify the $\tau^{\mbox{\tiny{ATE}}}$ from the observed data as $\tau^{\mbox{\tiny{ATE}}}=\operatorname{\mathbb{E}}\{Y_{i}Z_{i}/e(X_{i})\}-\operatorname{\mathbb{E}}\{Y_{i}(1-Z_{i})/(1-e(X_{i}))\}$. Among the consistent estimators of $\tau^{\mbox{\tiny{ATE}}}$, we focus on the Hájek inverse probability weighting estimator for analytical simplicity:
@@
Lunceford and Davidian, (2004) show that $N^{1/2}(\widehat{\tau}_{N}-\tau^{\mbox{\tiny{ATE}}})$ is asymptotically normal with variance
@@
In observational studies, the true propensity score is unknown and is replaced by an estimated one. Lunceford and Davidian, (2004) show that the asymptotic variance of the Hájek estimator with estimated propensity score, $V_{0}$, is $V$ subtracted by a complex quadratic term and therefore smaller. Calculation of the quadratic term usually requires individual data, which is unavailable in most sample size calculation. Therefore, we will proceed with $V$ henceforth, noting that it leads to a conservative (i.e. larger) sample size estimate than that based on $V_{0}$. We will investigate the discrepancy between $V$ and $V_{0}$ using simulations in Section 4.
With a fixed value of $V$, a $(1-\alpha)$ confidence interval for $\widehat{\tau}_{N}$ is $\widehat{\tau}_{N}\pm z_{1-\alpha/2}V^{1/2}N^{-1/2}$, where $z_{1-\alpha/2}$ is the $(1-\alpha/2)$ quantile of the standard normal distribution. Without loss of generality, we consider an one-sided level-$\alpha$ test for the null hypothesis $H_{0}:\tau=0$ against the alternative $H_{a}:\tau>0$:
@@
Standard derivation shows the minimal sample size $N$ to reach a desired power $\beta$ for a level-$\alpha$ test (4) with a fixed variance $V$ is (the proof is relegated to Supplementary Material A.1):
$N={\tau^{-2}}{\left(z_{1-\alpha}{V}^{1/2}-z_{1-\beta}{V_{0}}^{1/2}\right)^{2}}\leq{\tau^{-2}}{V}\left(z_{1-\alpha}+z_{\beta}\right)^{2}.$ (5)
In a randomized experiment, the propensity score is fixed and known, equaling the treatment proportion $r$, so that ${V}$ simplifies to $\left[\operatorname{\mathbb{V}}\{Y(1)\}+\operatorname{\mathbb{V}}\{Y(0)\}\right]/\{2r(1-r)\}$. Moreover, $\Pr\{Y(z)\}=\Pr(Y\mid Z=z)$, and thus variance of the potential outcomes $Y(z)$ can be estimated by the sample variances of the outcome in arm $z$, $s_{z}^{2}$. Following the convention, we impose $s_{0}=s_{1}=s$ and define the effect size as $\widetilde{\tau}=\tau/s$. Then, to detect an effect size $\widetilde{\tau}$ in a randomized experiment using a one-sided level-$\alpha$ test (4) with power $\beta$, the minimal sample size is
$N=\frac{s_{1}^{2}+s_{0}^{2}}{r(1-r)\tau^{2}}\left(z_{1-\alpha}+z_{\beta}\right)^{2}=\frac{\left(z_{1-\alpha}+z_{\beta}\right)^{2}}{r(1-r)\widetilde{\tau}^{2}}.$ (6)
@@
In observational studies, power calculation is more complicated because (i) the propensity scores $e(X)$ are not known, (ii) $\Pr(Y(z))\neq\Pr(Y\mid Z=z)$ due to confounding, and thus the terms involving potential outcomes, $\operatorname{\mathbb{E}}\{Y(z)\}$ and $\operatorname{\mathbb{V}}\{Y(z)\}$, deviate from their observed counterparts $\operatorname{\mathbb{E}}(Y\mid Z=z)$ and $\operatorname{\mathbb{V}}(Y\mid Z=z)$, and (iii) the numerator and denominator inside the expectation operator in (3) can not be separated because the potential outcomes may be dependent of the propensity scores. When sampled units are available, we can readily address these complications. However, power calculation is conducted prior to data collection, with little information to estimate the propensity scores or calculate the expectations. Nonetheless, analysts usually have some summary information from historical studies about the intended population.
@@ 3 Sample size and power calculation
The essence of sample size calculation lies in computing the variance $V$ in (3). If we could observe $(X_{i},Z_{i},Y_{i})$ for each sampled unit $i$, $V$ could easily be estimated by sample moments, but not so with only summary data. To determine which summary information is sufficient for computing $V$, we rewrite the terms in expression (3) as (taking $z=1$ for example),
$\mathbb{E}\left[\frac{[Y(1)-\mathbb{E}\{Y(1)\}]^{2}}{e(X)}\right]=\mathbb{E}\left[{[Y(1)-\mathbb{E}\{Y(1)\}]^{2}}\right]\mathbb{E}\left\{\frac{1}{e(X)}\right\}+\mathrm{Cov}\left[{[Y(1)-\mathbb{E}\{Y(1)\}]^{2}},\frac{1}{e(X)}\right].$ (7)
This expression suggests that we need three components: (i) mean and variance of the potential outcome in each arm, (ii) mean of (the inverse of) the propensity score, and (iii) covariance between the potential outcomes and propensity score. We study each component in detail below.
@@ 3.1 Normal approximation of linear combinations of covariates
Both the propensity score and the outcome models are commonly specified as a generalized linear model with a linear combination of the covariates, $X^{\prime}\beta=\sum_{j=1}^{p}\beta_{j}X_{j}$ as the predictor. However, in power calculation, we usually do not have individual data or even summary statistics of each covariate, without which we cannot characterize the distributions of the propensity score or the outcome. Previous authors often operated with a single normal predictor (e.g. Raudenbush,, 1997; Kelcey et al.,, 2017; Qin,, 2024), but without theoretical justification. Here we present a special case of the Lyapunov Central Limit Theorem (Billingsley,, 1995) to approximate the distribution of the scalar summary $X^{\prime}\beta$ and thus justify such a procedure.
@@
The proof of Theorem 1 is relegated to Supplementary Material A.2. Condition (i) ensures that the distribution of each covariate does not have heavy tails, and condition (ii) ensures that no single covariate dominates the linear combination. Though these two conditions may not always hold, empirical evidence from simulated and three real applications (see Supplementary Material B.1) shows the normality generally holds in a wide range of scenarios.

### v4

@@ Abstract
This paper investigates the theoretical foundation and develops analytical formulas for sample size and power calculations for causal inference with observational data. By analyzing the variance of an inverse probability weighting estimator of the average treatment effect, we decompose the power calculation into three components: propensity score distribution, potential outcome distribution, and their correlation. We show that to determine the minimal sample size of an observational study, in addition to the standard inputs in the power calculation of randomized trials, it is sufficient to have two parameters, which quantify the strength of the confounder-treatment and the confounder-outcome association, respectively. For the former, we propose using the Bhattacharyya coefficient, which measures the covariate overlap and, together with the treatment proportion, leads to a uniquely identifiable and easily computable propensity score distribution. For the latter, we propose a sensitivity parameter bounded by the R-squared statistic of the regression of the outcome on covariates. Our procedure relies on a parametric propensity score model and a semiparametric restricted mean outcome model, but does not require distributional assumptions on the multivariate covariates. We develop an associated R package PSpower.
@@ 1 Introduction
Randomized controlled trials are the gold standard for causal inference. Designing randomized trials requires sample size and power calculations, in which investigators pre-specify a target estimand—typically the effect size of a treatment—and characteristics of the sample of the intended study. The goal is to determine the minimal sample size needed to achieve the statistical power ($\beta$), given a type I error rate ($\alpha$) level, to test the treatment effect estimated by a specific method based on a sample that satisfies these characteristics, or conversely, to determine the power to detect a treatment effect at the $\alpha$ level given a fixed sample size. Sample size and power calculations are the two sides of the same coin, and we will use them interchangeably henceforth. Power calculation is conducted in the design stage of a study, prior to data collection. Thus, individual data are usually unavailable; investigators at best have some summary information from previous similar studies, e.g. effect sizes. Analysts must rely on this limited summary information to approximate the full data structure. The power calculation for randomized trials is largely simplified due to randomization and has been well established (Chow et al., 2017; Branson et al., 2024).
@@
When randomized trials are not feasible due to logistical or ethical constraints, researchers are increasingly using observational data to emulate randomized trials. The power analysis for observational studies is challenging and has been less developed. In randomized trials, there is no confounding due to randomization; therefore, one can simply use the difference in mean outcomes between arms to estimate treatment effects. Closed-form formulas, e.g., based on the two-sample z-test, are readily applicable for sample size calculation. In contrast, in observational studies, confounding must be adjusted for to obtain an unbiased estimate of treatment effects. Applying power calculation formulas designed for randomized trials to observational studies often leads to severely underpowered studies. In practice, a general power analysis method is to use simulations to generate individual data based on pre-specified summary statistics (Mukherjee et al., 2003; Qin, 2024). This method usually requires unverifiable assumptions of the entire data generating process and lacks a theoretical basis on what summary statistics are sufficient or necessary. In fact, many different data generating models may satisfy the same summary statistics but lead to vastly different sample sizes; choosing between these models is usually ad hoc. Another method is to use the variance inflation factor (Hsieh and Lavori, 2000; Shook-Sa and Hudgens, 2022) to approximate the inflation in the sample size of an observational study compared to a randomized trial. For example, Shook-Sa and Hudgens (2022) derived a variance inflation factor formula based on the analysis of the variance of an inverse probability weighting estimator of the average treatment effect; however, their formula requires knowledge on the distribution of the weights, which is usually not available.
@@
In summary, existing power calculation methods usually require information on the data generating process or the distributions of the weights or covariates, which is not always available in the design stage of a study. In this paper, we bypass this limitation by reducing the variance formula to a function of a few interpretable and commonly available summary parameters instead of the weights or covariates. Based on the analysis of variance of the Hájek inverse probability weighting estimator, we decompose the power calculation into three components: propensity score distribution, potential outcome distribution, and their correlation. We show that to determine the minimal sample size of an observational study, in addition to the standard inputs in the power calculation of a randomized trial, it is sufficient to have two parameters, which quantify the strength of the confounder-treatment and the confounder-outcome association, respectively. For the former, we propose using the Bhattacharyya coefficient, which measures the covariate overlap and, together with the treatment proportion, leads to a uniquely identifiable and easily computable propensity score distribution. For the latter, we propose a sensitivity parameter bounded by the R-squared statistic of the regression of the outcome on covariates. Our derivation relies on a parametric logistic probability score model and a semiparametric restricted mean outcome model with homogeneous treatment effect. However, by appealing to the Lyapunov-type Central Limit Theorem, we bypass specific distributional assumptions on the multivariate covariates. We derive analytical formulas for three common causal estimands: the average treatment effect for the overall (ATE), treated (ATT), and overlapped population (ATO). Simulations suggest that our formula is robust under treatment heterogeneity. We develop an associated R package PSpower, available at https://CRAN.R-project.org/package=PSpower.
@@ 2 Variance and sample size based on the Hájek estimator
Suppose we have a sample of $N$ units. Each unit $i$ ($i=1,2,\ldots,N$) has a binary treatment indicator $Z_{i}$, with $Z_{i}=0$ being the control and $Z_{i}=1$ being treated, and a vector of $p$ covariates $X_{i}=(X_{1i},\cdots,X_{pi})^{\prime}$. For each unit $i$, there are two potential outcomes $\{Y_{i}(1),Y_{i}(0)\}$ mapped to treatment and control status, of which only the one corresponding to the observed treatment is observed, denoted by $Y_{i}=Z_{i}Y_{i}(1)+(1-Z_{i})Y_{i}(0)$; the other potential outcome is missing. The propensity score is the probability of a unit being assigned to the treatment group given the covariates (Rosenbaum and Rubin, 1983): $e(x)=\Pr(Z_{i}=1\mid X_{i}=x)$.
The goal is to use the sample to estimate the average treatment effect (ATE):
@@
where $\tau(x)=\operatorname{\mathbb{E}}\{Y_{i}(1)-Y_{i}(0)\mid X_{i}=x\}$ is the conditional average treatment effect at covariate value $x$. To identify $\tau^{\mbox{\tiny{ATE}}}$, we assume SUTVA (Rubin, 1980) and maintain the two standard assumptions: (i) unconfoundedness, $\Pr(Z_{i}\mid Y_{i}(0),Y_{i}(1),X_{i})=\Pr(Z_{i}\mid X_{i})$; (ii) overlap, $0<\Pr(Z_{i}=1\mid X_{i})<1$. Among the consistent estimators of $\tau^{\mbox{\tiny{ATE}}}$, we focus on the Hájek inverse probability weighting estimator for analytical simplicity:
@@
Lunceford and Davidian (2004) show that $N^{1/2}(\widehat{\tau}_{N}-\tau^{\mbox{\tiny{ATE}}})$ is asymptotically normal with variance
@@
In observational studies, the true propensity score is unknown and is replaced by an estimated one. Lunceford and Davidian (2004) show that the asymptotic variance of the Hájek estimator with the propensity score estimated from a parametric model, $V_{0}$, is $V$ subtracted by a complex quadratic term and therefore smaller. Calculation of the quadratic term usually requires individual data, which is unavailable in power calculations. Therefore, we will proceed with $V$ henceforth, noting that it leads to a conservative (i.e. over-estimated) sample size estimate than that based on $V_{0}$. We investigate the discrepancy between $V$ and $V_{0}$ using simulations in Section 5. Besides the inverse probability weighting estimators, other two classes of estimators of $\tau^{\mbox{\tiny{ATE}}}$ are the outcome regression and doubly-robust estimators. However, asymptotic variances of these estimators depend on the specific outcome model, which is infeasible to pre-specify in the stage of power analysis. In contrast, as shown later, the Hájek estimator requires much weaker assumptions on the outcome model.
With a fixed value of $V$, a $(1-\alpha)$ confidence interval for $\widehat{\tau}_{N}$ is $\widehat{\tau}_{N}\pm z_{1-\alpha/2}V^{1/2}N^{-1/2}$, where $z_{1-\alpha/2}$ is the $(1-\alpha/2)$ quantile of the standard normal distribution. Without loss of generality, we consider a one-sided level-$\alpha$ test for the null hypothesis $H_{0}:\tau=0$ against the alternative $H_{a}:\tau>0$:
@@
Standard derivation shows the minimal sample size $N$ to reach a desired power $\beta$ for a level-$\alpha$ test (4) with a fixed variance $V$ is (the proof is relegated to Appendix A.1):
$N={\tau^{-2}}{\left(z_{1-\alpha}{V}^{1/2}-z_{1-\beta}{V_{0}}^{1/2}\right)^{2}}\leq{\tau^{-2}}{V}\left(z_{1-\alpha}+z_{\beta}\right)^{2},$ (5)
assuming $\alpha<0.5$ and $\beta>0.5$, which commonly holds in practice.
In a randomized experiment, the propensity score is fixed and known, equalling the treatment proportion $r$, so V{V} simplifies to $\operatorname{\mathbb{V}}\{Y(1)\}/r+\operatorname{\mathbb{V}}\{Y(0)\}/(1-r)$. Moreover, $\Pr\{Y(z)\}=\Pr(Y\mid Z=z)$, and thus the variance of the potential outcomes $Y(z)$ can be estimated by the sample variances of the outcome in arm $z$, $s_{z}^{2}$. Following the convention, we impose $s_{0}=s_{1}=s$ and define the effect size as $\widetilde{\tau}=\tau/s$. Then, to detect an effect size $\widetilde{\tau}$ in a randomized experiment using a one-sided level-$\alpha$ test (4) with power $\beta$, the minimal sample size is
$N=\left(\frac{s_{1}^{2}}{r}+\frac{s_{0}^{2}}{1-r}\right)\frac{\left(z_{1-\alpha}+z_{\beta}\right)^{2}}{\tau^{2}}=\frac{\left(z_{1-\alpha}+z_{\beta}\right)^{2}}{r(1-r)\widetilde{\tau}^{2}}.$ (6)
@@
In observational studies, power calculation is more complicated because (i) the propensity scores $e(X)$ are not known, (ii) $\Pr(Y(z))\neq\Pr(Y\mid Z=z)$ due to confounding, and thus the terms involving potential outcomes, $\operatorname{\mathbb{E}}\{Y(z)\}$ and $\operatorname{\mathbb{V}}\{Y(z)\}$, deviate from their observed counterparts $\operatorname{\mathbb{E}}(Y\mid Z=z)$ and $\operatorname{\mathbb{V}}(Y\mid Z=z)$, and (iii) the numerator and denominator inside the expectation operator in (3) can not be separated because the potential outcomes may be dependent on the propensity scores. When sampled units are available, we can readily address these complications. However, power calculation is conducted prior to data collection, with at best limited summary information to estimate the propensity scores or calculate the expectations.
@@ 3 Sample size and power calculation
The essence of power calculation lies in computing the variance $V$ in (3). If we could observe $(X_{i},Z_{i},Y_{i})$ for each sampled unit $i$, $V$ could easily be estimated by sample moments, but not so with only summary data. To determine which summary information is sufficient for computing $V$, we rewrite the terms in expression (3) as (taking $z=1$ for example),
$\mathbb{E}\left[\frac{[Y(1)-\mathbb{E}\{Y(1)\}]^{2}}{e(X)}\right]=\mathbb{E}\left[{[Y(1)-\mathbb{E}\{Y(1)\}]^{2}}\right]\mathbb{E}\left\{\frac{1}{e(X)}\right\}+\mathrm{Cov}\left[{[Y(1)-\mathbb{E}\{Y(1)\}]^{2}},\frac{1}{e(X)}\right].$
This expression suggests that we need three components: the propensity score distribution, the potential outcome distribution, and their correlation. We study each component in detail below.
@@ 3.1 Normal approximation of linear combinations of covariates
Both the propensity score and the outcome models are commonly specified as a generalized linear model with a linear combination of covariates, $X^{\prime}\beta=\sum_{j=1}^{p}\beta_{j}X_{j}$, as the predictor. However, in power calculations, we usually do not have specific information on each covariate. Previous authors often operated with a single normal predictor (e.g. Raudenbush, 1997; Kelcey et al., 2017), but without theoretical justification. Here we invoke a special case of the Lyapunov Central Limit Theorem (Billingsley, 1995) to approximate the distribution of the scalar summary $X^{\prime}\beta$ and thus justify such a procedure.
@@
The proof is relegated to Appendix A.2. Condition (i) ensures that the distribution of each covariate does not have heavy tails, and condition (ii) ensures that no single covariate dominates the linear combination. Theorem 1 suggests that, under Condition (i)-(ii), the linear combination $X^{\prime}\beta$ converges to a Normal distribution as the number of covariates ($p$) increases. In practice, some covariates might be correlated and $p$ might be modest. Some versions of the central limit theorem for dependent variables can be invoked when there is known correlation structure between covariates. However, in the context of power analysis, such information is usually not available and, in fact, analysts would avoid making specific assumptions about the covariate distributions. Instead, we conducted extensive simulations to examine the empirical distribution of $X^{\prime}\beta$ under a wide range of scenarios. These simulations show $X^{\prime}\beta$ approximates a Normal distribution satisfactorily often with as few as 10 variables as long as there is no strong pairwise correlation between $X$. We also draw the empirical distribution of $X^{\prime}\beta$ in four real applications, including the benchmark RHC data used in Section 5.3, with $p$ between 10 to 50, each of which also exhibits normality (details in Appendix B.1). Importantly, our derivation in the following sections only relies on the normality of $X^{\prime}\beta$ rather than the exact distribution, which is easier to achieve with a modest $p$ and mild correlation between covariates.